Please , I need full explanation and full answers .

Thank youThe enthalpy of solution of KBr in water is about
Answer #1

Problem-2: Given that when the atmospheric pressure is 1 atm, the concentration of radon in water is 7.27x10-3 M

Now applying Henry's law

P = KH x C

=>  KH = P / C = 1 atm / 7.27x10-3 M = 137.6 atm / M

Now when P = 36 atm, the molar concentration of radon in water can be calculated as

P = KH x C

=> C = P /KH = 36 atm / 137.6 atm / M = 0.262 M (answer)

Problem-3:(a): Given molality of the solution, m = 0.0520 m

i.e 0.0520 mol of caffeine is present in 1000 g of solvent chloroform.

molar mass of caffeine = 194.19 g/mol

Hence mass of caffeine present in 1000 g of solvent chloroform = 0.0520 mol x 194.19 g/mol = 10.098 g

Hence percentage of caffaine by mass in the solution = [10.098 g / 1000 g] x 100 = 1.01 % (answer)

(b): molar mass of chloroform = 119.3776 g/mol

mass of chloroform = 1000 g

Hence moles of solvent chloroform = mass / molar mass = 1000 g / 119.3776 g/mol = 8.377 mol

mole fraction caffeine in the solution = moles of caffeine / (moles of caffeine + moles of chloroform)

= 0.0520 mol / (0.0520 mol + 8.377 mol)

= 0.00617 (answer)